OpenStudy (anonymous):

simplify by removing factors of 1 80^4y^6 /50v^2y^4 how do you do these now?

OpenStudy (anonymous):

180^4y^6/50v^2y^4 =>180^4y^2/50v^2

OpenStudy (anonymous):

Assuming the equation as \[\frac{180^{4} \times y^6}{50 v^2 \times y^4}\] ,then the Greatest common denominator of \[(180^{4} \times y^6)\] and \[(50 v^2 \times y^4)\] is \[\frac{18^4}{5\times y^4}\] so dividing numerator and denominator by this gcd we would get the simplified form \[\frac{18^4 \times y^2}{50\times v^2}\]